You need to write down all necessary steps/formulas leading to your results in EXCEL FORMAT:

A share of common stock just paid a dividend of $2.00. If the expected long-run growth rate for this stock is 5.40%, and if investors' required rate of return is 13.50%, what is the stock price?
Answer #1

D1 DO*(1+g) = 2*(1+0.054) = 2.108 Value of Stock Next year dividend/(Discount rate-Growth rate) 2.108/(0.135-0.054) $26.02